(25 PTS) 4. Use matrix method to find the Find the solution to the (IVP). O=-376) :01-62
Answer #1

Xglo) Eiganwalues of A : det (4-AI) - 0 -3-1 (I+A) (3+A) +2=0 d= -2-i,-2ti Thins Eigenvector of d= -2ti : [A-(-2ri)]x = o 1-iaigenvector For two distinct complex eigenv alues di tdz where M= qriß and dz=a-iß and corvesponding eigenvedor n2=u-iv, gane